A monopolist can produce at a constant average (and marginal) cost of AC = MC = $5.
It faces a market demand curve given by Q = 53 – P.
a. Calculate the profit-maximizing price and quantity for this monopolist. Also
calculate its profits.
1
Expert's answer
2020-09-24T14:37:27-0400

Since Q = 53 - P,

P = 53 - Q

Profit is maximized when MR = MC.

Total revenue (TR) = P1"\times" "Q_{1}" = 53Q - Q2

P1​= 53Q - Q2

MR = "\frac{dtr}{dq}"=53 - 2Q

53 - 2Q = 5

2Q = 48

Q = 24

P = 53 - 24 = 29

Profit = Q"\times"(P - AC) = 24"\times"(29 - 5) = 576